Find two consecutive integers such that 4 times the larger exceeds 3 times the smaller by 23. Steps would be helpful.
Answer by oscargut(686)

Find two consecutive integers such that 4 times the larger exceeds 3 times the smaller by 23. Steps would be helpful.
1st integer = n (smaller)
2nd integer = n+1 (larger)
4 times the larger exceeds 3 times the smaller by 23
then 4(n+1)=3n+23
then 4n+4=3n+23 then 4n-3n=23-4 then n=19 and n+1=20
R: 19 and 20